OpenStudy (anonymous):

what is the density of hydrogen sulfide (h2s) at 1.3 atm and 329 K?

OpenStudy (jfraser):

the ideal gas law can be manipulated to solve for density\[P*V = n*R*T\] since\[n = \frac{mass}{MM}\] then \[P*V = (\frac{g}{MM})*R*T\] the density of a substance is the ratio of mass/volume, so rearrange those pieces to solve for g/V \[\frac{g}{V} = \frac{MM*P}{R*T}\] Plug and chug.